# Simplified Mamba with Disentangled Dependency Encoding for Long-Term Time Series Forecasting

### v2

@@ 6 Experiments
We comprehensively evaluate the performance and efficiency of the proposed SAMBA in LTSF and analyze the effectiveness of each component. Meanwhile, we also extend our disentangled strategy to other models, validating the universality of the proposed framework. We briefly list the datasets and baselines below. More details are provided in Appendix B.
@@
Datasets. We conduct experiments on nine real-world datasets following Zhou et al. (2021) and Liu et al. (2023): (1) ECL, (2) ETTh1, (3) ETTh2, (4) ETTm1, (5) ETTm2, (6) Exchange, (7) Traffic, (8) Weather, (9) Solar-Energy.
@@
Baselines. We carefully use 13 popular LTSF forecasting models as our baselines and we cite their performance from Liu et al. (2023) if applicable. Our baselines include (1) Pretrained Language model-based model: FTP (Zhou et al. (2023)); (2) Transformer-based models: Autoformer (Wu et al. (2021)), FEDformer (Zhou et al. (2022)), Stationary (Liu et al. (2022b)), Crossformer (Zhang & Yan (2022)), PatchTST (Nie et al. (2022)), iTransformer (Liu et al. (2023)), CARD (Wang et al. (2024)); (3) Linear-based models: DLinear (Zeng et al. (2023)), TiDE (Das et al. (2023)), RLinear (Li et al. (2023)); and (4) TCN-based models: SCINet (Liu et al. (2022a)), TimesNet (Wu et al. (2022)).
6.1 Long-Term Time Series Forecasting
@@
Table 4 presents the multivariate long-term forecasting results. In general, SAMBA outperforms all baseline methods. Specifically, for the ETT dataset, which exhibits weak cross-variate dependency, models using the CD strategy, such as iTransformer and Crossformer, underperform the models that use the CI strategy (e.g. PatchTST). However, SAMBA, which introduces cross-variate dependency through a disentangled encoding strategy, demonstrates superior performance over all models that use the CI strategy. For datasets with significant cross-variate dependency, such as Weather, ECL, and Traffic, SAMBA also performs comparable to or superior to the SOTA iTransformer.
@@
Table 4: Multivariate forecasting results with prediction lengths $S\in\{96,192,336,720\}$ and fixed lookback length $T=96$. Results are averaged from all prediction lengths. Full results are listed in Appendix L.2.
@@ 6.2 Ablation Experiment
To verify the rationale for removing the non-linear activation function and the disentangled encoding strategy, we choose our proposed SAMBA as the SOTA benchmark and develop three variant models. Figure 4 shows that SAMBA consistently outperforms other variants. Excluding either the nonlinear activation function or any branch of the disentangled encoding strategy results in a performance decline, demonstrating the effectiveness of our architecture and strategy. After incorporating the patch technique, the impact of the nonlinear activation function decreased. This occurs because the patch technique enriches the information available for each token, thereby enhancing the capture of semantic dependency. Additionally, our analysis shows that removing the cross-variate dependency encoding branch leads to a more significant performance degradation on the Traffic dataset compared to other datasets. This result is intuitive, as variates in traffic scenarios typically exhibit strong interdependencies (Han et al. (2024)).
@@
Figure 4: Ablation study of removing nonlinear activation function and disentangled encoding strategy in SAMBA. 4 cases are include: 1) SAMBA; 2) SAMBA without time encoding; 3) SAMBA without variate encoding; 4) Mamba with disentangled encoding.
6.3 Evaluating the Universality of Disentangled Encoding Strategy
@@
We utilize Transformer and their variants, including Informer (Zhou et al. (2021))and PatchTST (Nie et al. (2022)), to validate the universality of our proposed disentangled encoding strategy. Table 6 has demonstrated our strategy is effective not only for Mamba but also for Transformer-based models ( Full results can be found in Appendix L.1.). The significant improvements indicate that these models have not properly modeled and introduced cross-variate dependency, whereas our strategy effectively mitigates these issues. Therefore, this disentangled encoding strategy is a model-agnostic method and could potentially be integrated with other models.
6.4 Efficiency Analysis of SAMBA
In general, the time complexity of SAMBA is $O(D^{2}\cdot NJ+D\cdot NJ\log(NJ))$ where $D$ is the dimension of token embedding, $T$ is the length of tokens, $N$ is the number of variates, $J$ is the number of patches ( The detailed analysis process is given in Appendix D). This validates that SAMBA maintains the near linear time complexity w.r.t. of the patch number and the variate number, a key merit of Mamba, while also inheriting the efficiency of the patching strategy.
@@ 6.4 Efficiency Analysis of SAMBA
In addition, we provide a detailed efficiency comparison between SAMBA and the baseline models. As shown in Table 6, SAMBA achieves both a faster training speed and a smaller memory usage compared to many SOTA transformer-based models, such as PatchTST and Crossformer, which also employ attention mechanisms in temporal dimensions. Furthermore, SAMBA exhibits a much smaller increase in memory consumption and training time as input lengths grow, underscoring its superior overall efficiency.
@@
Table 5: Performance gains of our disentangled encoding strategy on MAE and MSE.
@@
SAMBA 448 0.0096 684 0.0145 1028 0.0268
PatchTST 790 0.0095 1980 0.0222 3264 0.0398
iTransformer 630 0.0113 632 0.0116 634 0.0121
DLinear 338 0.0043 342 0.0049 344 0.0055
TimesNet 930 0.0653 1446 0.1099 2212 0.243
Crossformer 1736 0.0502 2304 0.0664 3322 0.1031
FEDFormer 2336 0.2021 2772 0.2262 3616 0.2365
Autoformer 2502 0.0517 4408 0.0982 7922 0.1716
Table 6: Efficiency Analysis: The GPU memory (MiB) and speed (running time, s/iter) of each model.
7 Conclusions and Future Work
In this paper, we identify and rigorously define three crucial dependencies in time series: order dependency, semantic dependency, and cross-variate dependency. Based on these dependencies, we examine the three limitations of existing time series forecasting models: (1) The inability to capture order and semantic dependencies simultaneously. (2) The presence of overfitting issues. (3) The improper introduction of cross-variate dependency. We further empirically verify the superiority of Mamba as an alternative to linear and Transformer models, despite its redundant nonlinear activation functions. Building on these insights, we introduce SAMBA, a model derived from a simplified Mamba with a theoretically guaranteed disentangled encoding strategy. Experimentally, SAMBA achieves SOTA performance, validating the rationality and efficacy of our method design grounded in rich empirical findings. Our efforts further demonstrate the great potential of Mamba in time series modeling. In the future, we aim to delve deeper into diverse time series tasks, pushing the boundaries of analysis with Mamba.
@@ A.1 Proof for Theorem 1
$\mathcal{F}_{d}(\mathbf{E})^{1}$ $=\text{FFN}\left(\Phi(\mathbf{E}^{1})||\Psi\left(\mathbf{E}^{1},\mathbf{E}^{2}\right)^{1}\right),$ (5)
$\mathcal{F}_{ttv}(\mathbf{E})^{1}$ $=\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\Phi(\mathbf{E}^{2})\right)^{1}\right).$ (6)
@@
$\begin{split}H(\mathbf{c}^{1}|\mathcal{F}_{ttv}(\mathbf{E})^{1})&=H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\Phi(\mathbf{E}^{2})\right)^{1}\right)\right)\\ &\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\Phi(\mathbf{E}^{2})\right)^{1}\right)\right)\\ &=H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\phi(\mathcal{O}(\mathbf{E}^{2}),\mathcal{S}(\mathbf{E}^{2}))+\mathcal{Z}(\mathbf{E}^{2})\right)^{1}\right)\right)\\ &\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right),\mathcal{S}(\mathbf{E}^{2}),\mathcal{Z}(\mathbf{E}^{2})\right).\end{split}$ (7)
@@
$H(\mathbf{c}^{1}|\mathcal{F}_{ttv}(\mathbf{E})^{1})\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right),\mathcal{S}(\mathbf{E}^{2})\right).$ (8)
@@
$\begin{split}H(\mathbf{c}^{1}|\mathcal{F}_{ttv}(\mathbf{E})^{1})&\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right),\mathcal{S}(\mathbf{E}^{2})\right)\\ &=H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right)\right)\\ &-I\left(\mathbf{c}^{1};\mathcal{S}(\mathbf{E}^{2})|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right)\right)\\ &=H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right)\right)-I(\mathbf{c}^{1};\mathcal{S}(\mathbf{E}^{2})|\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2}))\\ &\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right)\right)-\epsilon\\ &\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Phi(\mathbf{E}^{1})||\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right)\right)-\epsilon\\ &\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Phi(\mathbf{E}^{1})||\mathcal{O}(\mathbf{E}^{2})\right)\right)-\epsilon.\end{split}$ (9)
@@
For $\mathcal{F}_{d},$ though the cross-variate dependency encoding $\Psi(\cdot)$ does not explicitly model the order dependencies in $\mathbf{E}^{2},$ it aggregates the patch embedding of the 2nd variate to the target variate at each time step. This enables the $\text{FFN}(\cdot)$ layer, which transforms the input with linear mapping, to further capture the order dependencies. Moreover, according to the empirical findings in Table 1 as well as previous works Zeng et al. (2023); Li et al. (2023), we assume that linear mapping is not worse than other complex cross-time encoders such as Transformers and Mamba in terms of order dependency modeling. Therefore, combining inequality (9), we have
$H(Y_{1}|F_{d}(\mathbf{E})^{1})$ $=H(Y_{1}|\text{FFN}\left(\Phi(\mathbf{E}^{1})||\Psi\left(\mathbf{E}^{1},\mathbf{E}^{2}\right)^{1}\right))$ (10)
$=H\left(Y_{1}|\text{FFN}\left(\Phi(\mathbf{E}^{1})\right)||\text{FFN}\left(\Psi\left(\mathbf{E}^{1},\mathbf{E}^{2}\right)^{1}\right)\right)$ (11)
$\leq H(Y_{1}|\text{FFN}\left(\Phi(\mathbf{E}^{1})\right)||\mathcal{O}(\mathbf{E}^{2}))$ (12)
$\leq H(Y_{1}|\text{FFN}\left(\Phi(\mathbf{E}^{1})||\mathcal{O}(\mathbf{E}^{2})\right))$ (13)
$\leq H(Y_{1}|F_{ttv}(\mathbf{E})^{1})+\epsilon.$ (14)
A.2 Rationality of the Assumption (2) in Theorem 1

### v3

@@ 6. Experiments
We comprehensively evaluate the proposed SDE framework’s performance and efficiency in LTSF and analyze the effectiveness of each component. Meanwhile, we extend our disentangled strategy to other models, validating the proposed framework’s universality. We briefly list the datasets and baselines below. The implementation details are provided in Appendix 6.1.
@@
Datasets. We conduct experiments on nine real-world datasets following (Liu et al., 2023): (1) ECL, (2) ETTh1, (3) ETTh2, (4) ETTm1, (5) ETTm2, (6) Exchange, (7) Traffic, (8) Weather, (9) Solar-Energy.
@@
Baselines. We carefully use 14 popular LTSF forecasting models as our baselines and cite their performance from (Liu et al., 2023) if applicable. Our baselines include (1) Pretrained Language model-based model: FTP ((Zhou et al., 2023)); (2) Transformer-based models: Autoformer (Wu et al., 2021), FEDformer (Zhou et al., 2022), Stationary (Liu et al., 2022a), Crossformer (Zhang and Yan, 2022), PatchTST (Nie et al., 2022), iTransformer (Liu et al., 2023), CARD (Wang et al., 2024b); (3) Linear-based models: DLinear (Zeng et al., 2023), TiDE (Das et al., 2023), RLinear (Li et al., 2023); (5) RNN-based model: SegRNN (Lin et al., 2023); and (4) TCN-based models: SCINet (Liu et al., 2022b), TimesNet (Wu et al., 2022).
6.1. Experiment Details
All experiments are implemented using PyTorch on a single NVIDIA A40 GPU. We optimize the model using ADAM with an initial learning rate in the set $\{10^{-3},5\times 10^{-4},10^{-4}\}$ and L2 loss. The batch size is uniformly set to $\{32-128\}$, depending on the maximum GPU capacity. To ensure that all models can be fitted effectively, the number of training epochs is fixed at $10$ with an early stopping mechanism. By default, our SDE-Mamba encodes both the time and variate dimensions using a single-layer SDE-Mamba encoding, with a patch length of $16$ and a stride of $8$.
The baseline models we use for comparison are implemented based on the Time Series Library ((Wu et al., 2022)) repository, and they are run with the hyperparameter settings they recommend. For CARD, we download the model from https://github.com/wxie9/CARD/tree/main, but utilize the same training loss as the other models (i.e., the Time Series Library GitHub repository) recommend for a fair comparison. Additionally, the full results of the predictions come from the outcomes reported by the iTransformer ((Liu et al., 2023)), ensuring fairness.
6.2. Long-Term Time Series Forecasting
@@
Table 4 presents the multivariate long-term forecasting results. In general, SDE-Mamba outperforms all baseline methods. Specifically, for the ETT dataset, which exhibits weak cross-variate dependency, models using the CD strategy, such as iTransformer and Crossformer, underperform the models that use the CI strategy (e.g., PatchTST). However, SDE-Mamba, which introduces cross-variate dependency through a disentangled encoding strategy, demonstrates superior performance over all models that use the CI strategy. For datasets with significant cross-variate dependency, such as Weather, ECL, and Traffic, SDE-Mamba also performs comparable to or superior to the SOTA iTransformer. Furthermore, the table highlights a remarkable performance boost of SDE-SegRNN over its predecessor, SegRNN, reinforcing the efficacy of our SDE framework in improving state-space models (SSMs). For instance, on the Solar-Energy dataset, SDE-SegRNN reduces the MSE from 0.239 (SegRNN) to 0.225, while on Traffic, it improves the MSE from 0.670 (SegRNN) to 0.622, confirming its ability to enhance predictive accuracy across diverse datasets. These results validate the effectiveness of our proposed framework in advancing multivariate long-term forecasting.
@@
Table 4. Multivariate forecasting results with prediction lengths $S\in\{96,192,336,720\}$ and fixed lookback length $T=96$. Results are averaged from all prediction lengths. Full results are listed in Table 8.
@@ 6.3. Ablation Study
To verify the rationale for removing the non-linear activation function and the disentangled encoding strategy, we choose our proposed SDE-Mamba as the SOTA benchmark and develop three variant models. Figure 4 shows that SDE-Mamba consistently outperforms other variants. Excluding either the nonlinear activation function or any branch of the disentangled encoding strategy results in a performance decline, demonstrating the effectiveness of our architecture and strategy. After incorporating the patch technique, the impact of the nonlinear activation function decreased. This occurs because the patch technique enriches the information available for each token, thereby enhancing the capture of semantic dependency. Additionally, our analysis shows that removing the cross-variate dependency encoding branch leads to a more significant performance degradation on the Traffic dataset compared to other datasets. This result is intuitive, as variates in traffic scenarios typically exhibit strong interdependencies (Han et al., 2024).
@@
Figure 4. Ablation study of removing nonlinear activation function and disentangled encoding strategy in SDE-Mamba. 4 cases are include: 1) SDE-Mamba; 2) SDE-Mamba without time encoding; 3) SDE-Mamba without variate encoding; 4) Mamba with disentangled encoding.
6.4. Evaluating the Universality of Disentangled Encoding Strategy
@@
We utilize Transformer and their variants, including Informer (Zhou et al., 2021) and PatchTST (Nie et al., 2022), to validate the universality of our proposed disentangled encoding strategy. Table 7 has demonstrated our strategy is effective not only for Mamba but also for Transformer-based models. Significant improvements indicate that these models have not properly modeled and introduced cross-variate dependency, whereas our strategy effectively mitigates these issues. Therefore, this disentangled encoding strategy is a model-agnostic method and could potentially be integrated with other models.
6.5. Efficiency Analysis of SDE-Mamba
First, the time complexity of each Mamba block is $O(D^{2}T+DT\log T)$(Fu et al., 2022; Gu and Dao, 2023)), where $D$ is the dimension of token embeddings and $T$ is the length of tokens. As a SDE-Mamba block only removes the nonlinear activation before the selection mechanism, it possesses the same time complexity. Next, we consider the time complexity of the disentangled modeling. (i) The linear projection in the tokenization process takes time $O(NJPD).$ (ii) In the CI learning, the token length is the variate number $N,$ thereby taking time $O(N(D^{2}J+DJ\log J)).$ (iii) In the CD learning, the dimension of token embeddings becomes the patch number $J,$ while the token length becomes the hidden dimension $D.$ Considering the bidirectional design, it takes time $O(2J(D^{2}N+DN\log N)).$ (iv) The FFN-based aggregation operation and the final prediction head take time $O(NJD^{2}+NJDS).$ In general, the time complexity of SDE-Mamba is $O(NJPD+N(D^{2}J+DJ\log J)+2J(D^{2}N+DN\log N)+(NJD^{2}+NJDS))=O(D^{2}\cdot NJ+D\cdot NJ\log(NJ)).$ This validates that SDE-Mamba maintains the near linear time complexity w.r.t. of the patch number and the variate number, a key merit of Mamba, while also inheriting the efficiency of the patching strategy.
@@ 6.5. Efficiency Analysis of SDE-Mamba
In addition, we provide a detailed efficiency comparison between SDE-Mamba and the baseline models. As shown in Table 7, SDE-Mamba achieves both a faster training speed and a smaller memory usage compared to many SOTA transformer-based models, such as PatchTST and Crossformer, which also employ attention mechanisms in temporal dimensions. Furthermore, SDE-Mamba exhibits a much smaller increase in memory consumption and training time as input lengths grow, underscoring its superior overall efficiency.
6.6. Short-Term Forecasting
Table 5. Performance comparison of SDE-Mamba and ablation variants. 4 cases are include: 1) SDE-Mamba; 2) SDE-Mamba without time encoding; 3) SDE-Mamba without variate encoding; 4) Mamba with disentangled encoding.
SDE-Mamba w/o variat w/o time w/ af iTransformer
SMAPE 12.129 12.492 12.279 12.323 12.706
MASE 14.383 14.876 14.632 14.590 15.269
MAPE 1.652 1.697 1.661 1.693 1.737
OWA 0.878 0.904 0.887 0.897 0.923
In addition to long-term forecasting (LTSF), we also evaluate the proposed SDE-Mamba framework on a short-term forecasting task using the M4 dataset (Makridakis, 2018), which comprises 100,000 univariate time series from marketing domains sampled at yearly, quarterly, and monthly frequencies. As shown in Table 5, the full SDE-Mamba model consistently achieves the best performance across all four metrics, demonstrating its strong generalization ability beyond the LTSF setting. We further conduct ablation studies to assess the role of each component. All variants show performance drops in most metrics, highlighting the necessity of both time and variate encoding for modeling temporal and feature-specific patterns.
@@ 6.6. Short-Term Forecasting
Table 6. Performance gains of our disentangled encoding strategy on MAE and MSE.
@@
SDE-Mamba 2235 0.0403 2275 0.0711 2311 0.1232
PatchTST 3065 0.0658 12299 0.2382 25023 0.4845
iTransformer 3367 0.0456 3389 0.0465 3411 0.0482
DLinear 579 0.0057 619 0.0082 681 0.0139
TimesNet 6891 0.2492 7493 0.4059 9807 0.6289
Crossformer 21899 0.7941 40895 1.426 69771 2.424
FEDFormer 1951 0.1356 1957 0.1369 2339 0.1643
Autoformer 1489 0.0309 1817 0.0362 2799 0.0457
Table 7. Efficiency Analysis: The GPU memory (MiB) and speed (running time, s/iter) of each model on Traffic dataset. Mem means memory footprint.
6.7. Increasing Lookback Length
It is intuitive to expect that using a longer lookback window would enhance forecasting performance. However, previous research has shown that most transformer-based models cannot leverage the increased receptive field to improve predictive accuracy effectively ((Nie et al., 2022)). As a result, SOTA transformer-based models, such as iTransformer and PatchTST, have improved their design to more effectively utilize a longer lookback window. In evaluating SDE-Mamba, we explore two key questions: (1) Can SDE-Mamba effectively take advantage of a longer lookback window to achieve more accurate forecasting? (2) If so, how does its performance compare to other models that can also utilize a longer lookback window? For comparison, we select four baseline models—DLinear, PatchTST, and iTransformer. These models have been proven to utilize extended a lookback window effectively ((Liu et al., 2023; Nie et al., 2022)). As shown in Figure 5, we observe a steady decrease in MSE as the lookback window increases, confirming SDE-Mamba ’s capacity to learn from longer temporal contexts. Moreover, while both SDE-Mamba and the baseline models can effectively take advantage of a longer look-back window, SDE-Mamba consistently demonstrates superior overall performance.
Figure 5. Forecasting performance with the lookback length $T\in\{48,96,192,336,720\}$ and fixed prediction length $S=96$.
7. Conclusions and Future Work
In this paper, we identify and rigorously define three crucial dependencies in time series: order dependency, semantic dependency, and cross-variate dependency. Based on these dependencies, we examine the three limitations of existing time series forecasting models: (1) The inability to capture order and semantic dependencies simultaneously. (2) The presence of overfitting issues. (3) The improper introduction of cross-variate dependency. We further empirically verify the superiority of state space model as an alternative to linear and Transformer models, despite its redundant nonlinear activation functions. Building on these insights, we propose SDE, a novel framework designed to enhance the capability of SSMs for LTSF. Experimentally, SDE-Mamba achieves SOTA performance, validating the rationality and efficacy of our method design grounded in rich empirical findings. Our efforts further demonstrate the great potential of state space models in time series modeling. In the future, we aim to delve deeper into diverse time series tasks, pushing the boundaries of analysis with state space model.
Acknowledgements.
This work was supported by the National Key R&D Program of China (Grant No.2023YFF0725004), National Natural Science Foundation of China (Grant No.92370204), the Guangzhou Basic and Applied Basic Research Program under Grant No. 2024A04J3279, Education Bureau of Guangzhou Municipality, and CCF-DiDi GAIA Collaborative Research Funds.
@@ A.1. Proof for Theorem 1
(5) $\mathcal{F}_{d}(\mathbf{E})^{1}$ $=\text{FFN}\left(\Phi(\mathbf{E}^{1})||\Psi\left(\mathbf{E}^{1},\mathbf{E}^{2}\right)^{1}\right),$
(6) $\mathcal{F}_{ttv}(\mathbf{E})^{1}$ $=\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\Phi(\mathbf{E}^{2})\right)^{1}\right).$
@@
(7) $\begin{split}H(\mathbf{c}^{1}|\mathcal{F}_{ttv}(\mathbf{E})^{1})&=H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\Phi(\mathbf{E}^{2})\right)^{1}\right)\right)\\ &\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\Phi(\mathbf{E}^{2})\right)^{1}\right)\right)\\ &=H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\phi(\mathcal{O}(\mathbf{E}^{2}),\mathcal{S}(\mathbf{E}^{2}))+\mathcal{Z}(\mathbf{E}^{2})\right)^{1}\right)\right)\\ &\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right),\mathcal{S}(\mathbf{E}^{2}),\mathcal{Z}(\mathbf{E}^{2})\right).\end{split}$
@@
(8) $H(\mathbf{c}^{1}|\mathcal{F}_{ttv}(\mathbf{E})^{1})\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right),\mathcal{S}(\mathbf{E}^{2})\right).$
@@
(9) $\begin{split}H(\mathbf{c}^{1}|\mathcal{F}_{ttv}(\mathbf{E})^{1})&\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right),\mathcal{S}(\mathbf{E}^{2})\right)\\ &=H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right)\right)\\ &-I\left(\mathbf{c}^{1};\mathcal{S}(\mathbf{E}^{2})|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right)\right)\\ &=H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right)\right)\\ &-I(\mathbf{c}^{1};\mathcal{S}(\mathbf{E}^{2})|\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2}))\\ &\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right)\right)-\epsilon\\ &\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Phi(\mathbf{E}^{1})||\Psi\left(\Phi(\mathbf{E}^{1}),\mathcal{O}(\mathbf{E}^{2})\right)^{1}\right)\right)-\epsilon\\ &\geq H\left(\mathbf{c}^{1}|\text{FFN}\left(\Phi(\mathbf{E}^{1})||\mathcal{O}(\mathbf{E}^{2})\right)\right)-\epsilon.\end{split}$
@@
For $\mathcal{F}_{d},$ though the cross-variate dependency encoding $\Psi(\cdot)$ does not explicitly model the order dependencies in $\mathbf{E}^{2},$ it aggregates the patch embedding of the 2nd variate to the target variate at each time step. This enables the $\text{FFN}(\cdot)$ layer, which transforms the input with a linear mapping, to further capture the order dependencies. Moreover, according to the empirical findings in Table 3 as well as previous works (Zeng et al., 2023; Li et al., 2023), we assume that linear mapping is not worse than other complex cross-time encoders such as Transformers and Mamba in terms of order dependency modeling. Therefore, combining inequality (9), we have
(10) $H(Y_{1}|F_{d}(\mathbf{E})^{1})$ $=H(Y_{1}|\text{FFN}\left(\Phi(\mathbf{E}^{1})||\Psi\left(\mathbf{E}^{1},\mathbf{E}^{2}\right)^{1}\right))$
(11) $=H\left(Y_{1}|\text{FFN}\left(\Phi(\mathbf{E}^{1})\right)||\text{FFN}\left(\Psi\left(\mathbf{E}^{1},\mathbf{E}^{2}\right)^{1}\right)\right)$
(12) $\leq H(Y_{1}|\text{FFN}\left(\Phi(\mathbf{E}^{1})\right)||\mathcal{O}(\mathbf{E}^{2}))$
(13) $\leq H(Y_{1}|\text{FFN}\left(\Phi(\mathbf{E}^{1})||\mathcal{O}(\mathbf{E}^{2})\right))$
(14) $\leq H(Y_{1}|F_{ttv}(\mathbf{E})^{1})+\epsilon.$
A.2. Rationality of the Assumption (2) in Theorem 1